OpenStudy (anonymous):

help asap please? What is the electric force between a +2 µC point charge and a –2 µC point charge if they are separated by a distance of 5.0 cm? Show your work. (µC = 1.0 × 10–6 C)

OpenStudy (anonymous):

\[F = k\frac{Q_1Q_2}{r^2}=\frac{1}{4 \pi \epsilon_o}\frac{Q_1Q_2}{r^2}\] Be careful with your units!

OpenStudy (anonymous):

@Vandreigan , isn't it $$F = k\frac{\color{red}{|q_1||q_2|}}{r^2} $$ ?

OpenStudy (anonymous):

That's what I have written, but it isn't absolute values. The positive or negative sign of the result tells you if the force is attractive (the charges move together) or repulsive (charges move apart).